OpenStudy (nincompoop):

The area A of an equilateral triangle is a function of the length of one of its sides x. Find the formula for the given function and find the domain.

OpenStudy (anonymous):

so here, our independent variable is length of one of an equilateral triangle. Let x= length of one side, then height of the equilateral triangle would be h=xsin60. area is length * height * 1/2 A=x*(xsin60)*1/2= (x^2)sin60/2 =(sqrt3)x^2/4

OpenStudy (anonymous):

domain is that length of the equilateral sides is of real numbers, greater than 0.

OpenStudy (nincompoop):

why did you use a sin(60)?

OpenStudy (nincompoop):

I was thinking more along the lines of: A(x) = 1/2 b * h A(x) = 1/2 x * h A(x) = 1/2 x * sqrt[ x^2 - (x/2)^2] A(x) = 1/2 x * sqrt[x^2 - (x^2)/4] ...

OpenStudy (anonymous):

that is the same thing.

OpenStudy (anonymous):